OpenStudy (study_buddy99):

2 college alg problems? (equation down below)

OpenStudy (study_buddy99):

\[\frac{ 3x^2+11x-4 }{ 24x^3-8x^2 }\div \frac{ 9x+36 }{ 24x^4-36x^3 }\]

OpenStudy (study_buddy99):

find the product by factoring & reducing

OpenStudy (study_buddy99):

Find the sum \[\frac{ 1 }{ x^2 }+ \frac{ 1 }{ x^2+x }\]

rishavraj (rishavraj):

see \[3x^2 + 11x - 4 = (3x - 1)(x 4)\] similarly factorise and solve

OpenStudy (study_buddy99):

that makes no sense

OpenStudy (agent0smith):

Post one problem at a time, not two. For the first, start by factoring everything.

OpenStudy (study_buddy99):

\[\frac{ (3x-1)(x+4) }{ 8x^2 (3x-1)}\div \frac{ 6x^3(4x-6) }{ 9(x+4 }\] ?

OpenStudy (study_buddy99):

oops... at that point it would mult too

OpenStudy (agent0smith):

Looks about right. Now cancel off factors and simplify.

rishavraj (rishavraj):

@study_buddy99 u just need to factorize it and proceed :)

OpenStudy (study_buddy99):

I got\[\frac{ 6x^3(3x-1)(4x-6) }{ 8x^2(9) }\] is what I got... but that's not the answer

OpenStudy (agent0smith):

Keep simplifying...There's still common factors on top and bottom

rishavraj (rishavraj):

the first would look like : \[\frac{ (3x - 1)(x + 4) }{ 8x^2(3x - 1) }\] then u can cancel out like terms .....

rishavraj (rishavraj):

*first term

OpenStudy (agent0smith):

\[\large \frac{ (3x-1)(x+4) }{ 8x^2 (3x-1)}* \frac{ 6x^3(4x-6) }{ 9(x+4 )} = \frac{ 6x^3(4x-6) }{ 9*8x^2 } \]and you can still simplify...

OpenStudy (study_buddy99):

My teacher (and an answer checker both say it's \[\frac{ x(2x-3) }{ 6 }\]

OpenStudy (agent0smith):

Because you need to keep simplifying...\[\large \frac{ 6x^3(4x-6) }{ 9*8x^2 } = \frac{ 6x^3*2(2x-3) }{ 9*8x^2 } =\]

OpenStudy (study_buddy99):

okay, I see. I kept making a really stupid mistake... I got it now, thank you

OpenStudy (agent0smith):

6/9 is 2/3, 2/8 is 1/4\[\large \frac{ 6x^3*2(2x-3) }{ 9*8x^2 } =\frac{ 2x*1(2x-3) }{ 3*4 } =\]

OpenStudy (study_buddy99):

\[\frac{ 2x(4x+3)}{ 12 }\]

OpenStudy (agent0smith):

Is this a diff question...?

OpenStudy (agent0smith):

\[\large \frac{ 2x*1(2x-3) }{ 3*4 } = \frac{ x(2x-3) }{ 3*2 } \]

OpenStudy (study_buddy99):

no that's what I got.

OpenStudy (agent0smith):

You must've made some mistakes. For one, a negative sign became a positive. And a 2 became a 4.

OpenStudy (study_buddy99):

I'm stuck at \[\frac{ (6x^3)(2)(2x-3) }{ 8x^2(9) }\]

OpenStudy (study_buddy99):

But then where does the x^2 go and why can't I divide 6x^3/8x^2

OpenStudy (agent0smith):

x^3/x^2 = x 6/8 = 3/4 You can simplify the fractions however you like

OpenStudy (study_buddy99):

okay so I got it out to \[\frac{ 3x(2x+3) }{ 2(9) }\]

OpenStudy (study_buddy99):

then I simplified 3/18 to get \[\frac{ x(2x-3) }{ 6 }\]